Please, show clear work if you need thumbs up, be sure the answer is right, if the answer is incorrect you will get thumbs down.
1. One orange juice futures contract is on 15,000 pounds of frozen concentrate. Suppose that in September 2016 a company sells a March 2018 orange juice futures contract for 125 cents per pound. In December 2016, the futures price is 140 cents, in December 2017, it is 115 cents; and in February 2018, it is closed out at 130 cents. The company has a December year end a. b. c. What is the companys profit or loss on the contract? How is the profit or loss realized? What is the accounting and tax treatment of the transaction if the company is classified as (i) a hedger and (ii) a speculator?
Answer #1

The price goes up during the time the company holds the contract from 125 to 130 cents per pound. Overall the company therefore takes a loss of 15,000×0.05 = $750.

If the company is classified as a hedger this loss is realized in 2018,

If it is classified as a speculator it realizes

a loss of 15,000×0.15 = $2,250 in 2016,

a gain of 15,000×0.25 = $3,750 in 2017, and

a loss of 15,000×0.15 = $2,250 in 2018